# On local Galois deformation rings: generalised reductive groups

### v1

@@ 13. Complete intersection
In this section we will show that the deformation rings $R^{\square}_{G,\overline{\rho}}$ are complete intersection of expected dimension. As explained in the introduction it is enough to bound the dimension of their special fibre from above. We will achieve that by bounding the dimension of $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$.
@@ 13.1. Induction hypothesis
We will bound the dimension of $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ arguing by an induction that the assertion
@@
holds for all pairs $(H,\overline{\rho}_{1})$, where $H$ is a generalised reductive group scheme defined over a finite extension of $k$ and $\overline{\rho}_{1}:\Gamma_{F}\rightarrow H(\overline{k})$ is a continuous $H$-semisimple representation such that $\pi_{H}\circ\overline{\rho}_{1}:\Gamma_{F}\rightarrow(H/H^{0})(\overline{k})$ is surjective.
The induction argument takes place entirely in the special fibre and the group schemes $H$ appearing in the induction need not be special fibres of generalised group schemes defined over $\mathcal{O}$. We note that the schemes $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$ and $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$, $\overline{X}^{\mathrm{ps}}_{H}$ are still defined for such $H$ by replacing $\mathcal{O}$ with $k$ in all the definitions. All the statements proved for $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$, $\overline{X}^{\mathrm{ps}}_{G}$ when $G$ is a generalised reductive group scheme over $\mathcal{O}$, also hold for $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}$ and $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}$, $\overline{X}^{\mathrm{ps}}_{H}$. In particular, if $x$ is a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}_{1}}$ or a closed point of $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}_{1}}$ then the arguments of 5.16, 5.17 carry over and relate the completed local ring of $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$ at $x$ to the deformation ring $\overline{R}{}^{\square}_{\rho_{x},H}$, which parameterises deformations $\rho:\Gamma_{F}\rightarrow H(A)$ of $\rho_{x}$ for $A\in\mathfrak{A}_{\kappa(x)}$. If $H$ is a generalised reductive group scheme over $\mathcal{O}$ then $\overline{R}{}^{\square}_{\rho_{x},H}=R^{\square}_{\rho_{x},H}/\varpi$.
@@
This follows from [8, Lemma 3.5 (5)] applied to $\operatorname{Spec}R=\operatorname{Spec}S=Z$. ∎
@@
Let $Y_{\overline{\rho}^{\mathrm{ss}}}$ be the preimage in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ of the closed point in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and let $V$ be an open non-empty $G^{0}$-invariant subscheme of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$. Let $Z$ be the closure of $V$ in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Then $\dim Z=\dim V+1$.
@@
Since $V$ is $G^{0}$-invariant, its closure $Z$ is also $G^{0}$-invariant. It follows from the proof of [8, Lemma 3.21] that every irreducible component of $Z$ meets $Y_{\overline{\rho}^{\mathrm{ss}}}$ non-trivially. The assertion follows from [8, Lemma 3.18 (5)]. ∎
@@
From [8, Lemma 3.18] we deduce that
@@
We may assume that $\overline{X}^{\mathrm{git}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ is non-empty. Lemma 9.11 implies that $\dim L<\dim G$ and $\dim N_{k}\geq 1$. The assertion follows from Proposition 13.5. The assumption on $U$ implies that (58) changes to
(59) $\delta(U)\leq\dim L_{k}-\dim H.$
Thus (51) implies that
(60) $\begin{split}\dim G_{k}([F:\mathbb{Q}_{p}]+1)-&\dim Z\geq\\ &(\dim L_{k}-\dim H)([F:\mathbb{Q}_{p}]-1)+\dim N_{k}[F:\mathbb{Q}_{p}],\end{split}$
@@ 13.2. $W$-special locus
Thus $(\check{W}(A)/M)\otimes_{A}\kappa(x)$ is isomorphic to $\Gamma_{F}$-coinvariants of $\check{W}_{x}(-1)$. We conclude that $x\in X_{W}$ if and only if $\Gamma_{F}$-coinvariants of $\check{W}_{x}(-1)$ are non-zero, and hence $X_{W,0}=X_{W}$. It follows from upper-semicontinuity of the dimension function, [32, Example 12.7.2] that $X_{W,j}$ is a closed subscheme of $X_{W}$ for $j\geq 1$. ∎
@@
The proof is the same as the proof of [8, Lemma 3.21]. ∎
@@
The proof is the same as the proof of [8, Lemma 3.23]. ∎
@@
Since the $W$-special locus of level $j$ is closed in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ by Lemma 13.8 and $G^{0}$-invariant by 13.11, its image in $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is also closed by [52, Theorem 3 (iii)]. To prove part (2), let $\eta_{1},\ldots,\eta_{r}$ be the generic points of $C_{W,j}$. It is enough to find $H_{i}$ satisfying the required bounds on the dimension such that $\eta_{i}\in\overline{X}^{\mathrm{git}}_{H_{i}G,\overline{\rho}^{\mathrm{ss}}}$.
To ease the notation let $X$ be the $W$-special locus of level $j$ in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, let $\eta=\eta_{1}$ and let $\bar{\eta}$ be a geometric point above $\eta$. It follows from [52, Theorem 3] that there is $x\in X(\kappa(\bar{\eta}))$ above $\bar{\eta}$ such that the orbit $G^{0}\cdot x$ is closed in $X_{\bar{\eta}}$. Since $X$ is closed in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, Lemma 5.6 together with (28) imply that $X_{\bar{\eta}}$ is a closed $G^{0}$-invariant subscheme of $G^{N}_{\bar{\eta}}$, where $x\in X_{\bar{\eta}}$ is mapped to the $N$-tuple $y=(\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N}))$, where $\gamma_{i}$ are elements of $\Gamma_{F}$ chosen before Lemma 5.6. Hence, the orbit $G^{0}\cdot y$ is closed in $G^{N}_{\bar{\eta}}$ under the conjugation action. Let $H_{x}$ be the smallest closed subgroup of $G_{\bar{\eta}}$ containing the entries of $y$. It follows from [5, Theorem 3.1] that $H_{x}$ is a strongly reductive subgroup of $G_{\bar{\eta}}$, and thus is reductive by [40, Section 6]. It follows from Lemma 5.13 that $H_{x}$ is equal to the Zariski closure of $\rho_{x}(\Gamma_{F})$ in $G_{\bar{\eta}}$. Since $x\in X$, the dimension of the $H_{x}^{0}$-invariants of $W_{x}$ is at least $j+1$ by Lemma 13.10. If $\dim H_{x}=\dim G_{k}$ then $H_{x}^{0}=G_{\bar{\eta}}^{0}$ and hence $W^{G^{0}}\neq 0$. If $W^{G^{\prime}}=0$ then $\dim H_{x}\leq\dim G_{k}-2$ by 10.3.
By [40, Theorem 10.3] there is a closed reductive subgroup $H$ of $G$ defined over $\bar{k}$ and $g\in G(\kappa(\bar{\eta}))$ such that $H_{x}=g(H_{\bar{\eta}})g^{-1}$. Since $H$ is defined by finitely many polynomials with coefficients in $\bar{k}$, $H$ is defined over a finite extension of $k$. The point $x^{\prime}:=g(x)\in\overline{X}^{\mathrm{gen}}_{H}(\kappa(\bar{\eta}))$ maps to $\eta$ in $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$, thus $\eta$ lies in $\overline{X}^{\mathrm{git}}_{HG,\overline{\rho}^{\mathrm{ss}}}$. ∎
@@ 13.3. Bounding the dimension of the space
In this subsection we carry out the induction to prove (DIM). It follows from [43, Theorem 9.3 (3)] that (DIM) holds if $H^{0}$ is a torus. Thus we may assume that $\dim H\geq 3$.
@@
Let $\overline{{V}}_{P,ij}$ be the preimage of $\overline{U}_{P,ij}:=\overline{U}_{LG,\overline{\rho}^{\mathrm{ss}}}\cap\overline{X}^{\mathrm{git}}_{H_{i}L,\overline{\rho}^{\mathrm{ss}}}$ in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and let $\overline{Z}_{P,ij}$ be its closure. If part (a) holds then Corollary 13.6 gives us
@@
Lemma 9.11 implies that $\dim N_{k}\geq 1$, where $N$ is a unipotent radical of any R-parabolic with R-Levi $L$. It follows from Lemma 2.16, (73), (74) and (75) that $\dim\overline{Z}>\dim\overline{X}^{\mathrm{git}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ and thus $\dim\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}=\dim\overline{Z}$, which implies the last equality in (70). ∎
@@
If $\pi_{1}(G^{\prime})$ is étale or $x\in X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ then $(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})_{x}=(\operatorname{Lie}G^{\prime})_{x}$. It follows from Proposition 9.7 and Lemma 13.8 that $x\in V^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ if and only if $\rho_{x}$ is absolutely $G$-irreducible and $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))=0$. Since $x$ is closed in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$ its residue field is a local field by [8, Lemmas 3.17, 3.18]. Local Tate duality gives us an isomorphism $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))\cong H^{2}(\Gamma_{F},\operatorname{ad}^{0}\rho_{x})^{*}$, which implies the assertion. ∎
@@
where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map and $r=\dim G^{\prime}_{k}([F:\mathbb{Q}_{p}]+1)$. It is proved in [43, Corollary 9.5] that $X^{\mathrm{gen}}_{G/G^{\prime},\varphi\circ\overline{\rho}^{\mathrm{ss}}}$ is $\mathcal{O}$-flat of relative dimension $\dim(G/G^{\prime})_{k}([F:\mathbb{Q}_{p}]+1)$. It follows from 5.16 that $R^{\square}_{G/G^{\prime},\varphi\circ\rho_{x}}$ is flat over the coefficient ring $\Lambda$ for $\kappa(x)$ of relative dimension $\dim(G/G^{\prime})_{k}([F:\mathbb{Q}_{p}]+1)$. Thus $R^{\square}_{G,\rho_{x}}$ is flat over $\Lambda$ of relative dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$.
@@
For the purpose of establishing (DIM), after extending $\mathcal{O}$ by a finite extension we may assume, that $G^{0}$ is split, see 5.9. Let $\mu:=Z(G^{0})\cap G^{\prime}=Z(G^{\prime})$, $G_{1}:=G/\mu$ and let $\varphi_{1}:G\to G_{1}$ be the projection map. Since $Z(G^{0})$ is a split torus, $\mu$ is a finite diagonalisable group scheme and we may decompose $\mu$ canonically into a $p$-part and a prime-to-$p$ part, which are both normal in $G$. Let $x$ be a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}}$. 3.12 and 3.16 imply that $\dim\overline{R}^{\square}_{G,\rho_{x}}=\dim\overline{R}^{\square}_{G_{1},\varphi_{1}\circ\rho_{x}}$.

### v2

@@ 13. Complete intersection
In this section we will show that the deformation rings $R^{\square}_{G,\overline{\rho}}$ are complete intersections of expected dimension. As explained in the introduction it is enough to bound the dimension of their special fibres from above. We will achieve this by bounding the dimension of $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$.
@@ 13.1. Induction hypothesis
We will bound the dimension of $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ arguing by induction that the assertion
@@
holds for all pairs $(H,\overline{\rho}_{1})$, where $H$ is a generalised reductive group scheme defined over a finite extension of $k$ and $\overline{\rho}_{1}:\Gamma_{F}\rightarrow H(\overline{k})$ is a continuous $H$-semisimple representation such that $\pi_{H}\circ\overline{\rho}_{1}:\Gamma_{F}\rightarrow(H/H^{0})(\overline{k})$ is surjective. We refer the reader to Section 1.4.2 for the sketch of the proof strategy. Instead of reading this section linearly we suggest to first look at 13.25, where the induction step is carried out, and then look up the statements needed for its proof.
The induction argument takes place entirely in the special fibre and the group schemes $H$ appearing in the induction need not be special fibres of generalised reductive group schemes defined over $\mathcal{O}$. We note that the schemes $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$, $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$ and $\overline{X}^{\mathrm{ps}}_{H}$ are still defined for such $H$ by replacing $\mathcal{O}$ with $k$ in all the definitions. All the statements proved for $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $\overline{X}^{\mathrm{ps}}_{G}$ when $G$ is a generalised reductive group scheme over $\mathcal{O}$, also hold for $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}$, $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}$ and $\overline{X}^{\mathrm{ps}}_{H}$. In particular, if $x$ is a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}_{1}}$ or a closed point of $\overline{X}^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}_{1}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}_{1}}$ then the arguments of 5.16, 5.17 carry over and relate the completed local ring of $\overline{X}^{\mathrm{git}}_{H,\overline{\rho}_{1}^{\mathrm{ss}}}$ at $x$ to the deformation ring $\overline{R}{}^{\square}_{\rho_{x},H}$, which parameterises deformations $\rho:\Gamma_{F}\rightarrow H(A)$ of $\rho_{x}$ for $A\in\mathfrak{A}_{\kappa(x)}$. If $H$ is a generalised reductive group scheme over $\mathcal{O}$ then $\overline{R}{}^{\square}_{\rho_{x},H}=R^{\square}_{\rho_{x},H}/\varpi$.
@@
This follows from [6, Lemma 3.5 (5)] applied to $\operatorname{Spec}R=\operatorname{Spec}S=Z$. ∎
@@
Let $Y_{\overline{\rho}^{\mathrm{ss}}}$ be the preimage in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ of the closed point in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and let $V$ be an open non-empty $G^{0}$-invariant subscheme of $X\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$, where $X$ is either $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ or $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Let $Z$ be the closure of $V$ in $X$. Then $\dim Z=\dim V+1$.
@@
Since $V$ is $G^{0}$-invariant, its closure $Z$ is also $G^{0}$-invariant. It follows from the proof of [6, Lemma 3.21] that every irreducible component of $Z$ meets $Y_{\overline{\rho}^{\mathrm{ss}}}$ non-trivially. The assertion follows from [6, Lemma 3.18 (5)]. ∎
@@
From [6, Lemma 3.18] we deduce that
@@
We may assume that $\overline{X}^{\mathrm{git}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ is non-empty. Lemma 9.11 implies that $\dim L<\dim G$ and $\dim N_{k}\geq 1$. As in the proof of 13.6 the assertion follows from Proposition 13.5, but using (59) instead of (58):
(60) $\begin{split}\dim G_{k}([F:\mathbb{Q}_{p}]+1)-&\dim Z\geq\\ &(\dim L_{k}-\dim H)([F:\mathbb{Q}_{p}]-1)+\dim N_{k}[F:\mathbb{Q}_{p}].\end{split}$
@@ 13.2. $W$-special locus
Thus $(\check{W}(A)/M)\otimes_{A}\kappa(x)$ is isomorphic to $\Gamma_{F}$-coinvariants of $\check{W}_{x}(-1)$. We conclude that $x\in X_{W}$ if and only if $\Gamma_{F}$-coinvariants of $\check{W}_{x}(-1)$ are non-zero, and hence $X_{W,0}=X_{W}$. It follows from upper-semicontinuity of the dimension function, [33, Example III.12.7.2] that $X_{W,j}$ is a closed subscheme of $X_{W}$ for $j\geq 1$. ∎
@@
The proof is the same as the proof of [6, Lemma 3.21]. ∎
@@
The proof is the same as the proof of [6, Lemma 3.23]. ∎
@@
Since the $W$-special locus of level $j$ is closed in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ by Lemma 13.8 and $G^{0}$-invariant by 13.11, its image in $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is also closed by [57, Theorem 3 (iii)]. To prove part (2), let $\eta_{1},\ldots,\eta_{r}$ be the generic points of $C_{W,j}$. It is enough to find $H_{i}$ satisfying the required bounds on the dimension such that $\eta_{i}\in\overline{X}^{\mathrm{git}}_{H_{i}G,\overline{\rho}^{\mathrm{ss}}}$.
To ease the notation let $X$ be the $W$-special locus of level $j$ in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, let $\eta=\eta_{1}$ and let $\bar{\eta}$ be a geometric point above $\eta$. It follows from [57, Theorem 3] that there is $x\in X(\kappa(\bar{\eta}))$ above $\bar{\eta}$ such that the orbit $G^{0}\cdot x$ is closed in $X_{\bar{\eta}}$. Since $X$ is closed in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, Lemma 5.6 together with (28) imply that $X_{\bar{\eta}}$ is a closed $G^{0}$-invariant subscheme of $G^{N}_{\bar{\eta}}$, where $x\in X_{\bar{\eta}}$ is mapped to the $N$-tuple $y=(\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N}))$, where $\gamma_{i}$ are elements of $\Gamma_{F}$ chosen before Lemma 5.6. Hence, the orbit $G^{0}\cdot y$ is closed in $G^{N}_{\bar{\eta}}$ under the conjugation action. Let $H_{x}$ be the smallest closed subgroup of $G_{\bar{\eta}}$ containing the entries of $y$. It follows from [3, Theorem 3.1] that $H_{x}$ is a strongly reductive subgroup of $G_{\bar{\eta}}$, and thus is reductive by [42, Section 6]. It follows from Lemma 5.13 that $H_{x}$ is equal to the Zariski closure of $\rho_{x}(\Gamma_{F})$ in $G_{\bar{\eta}}$. Since $x\in X$, the dimension of the $H_{x}^{0}$-invariants of $W_{x}$ is at least $j+1$ by Lemma 13.10. If $\dim H_{x}=\dim G_{k}$ then $H_{x}^{0}=G_{\bar{\eta}}^{0}$ and hence $W^{G^{0}}\neq 0$. If $W^{G^{\prime}}=0$ then $\dim H_{x}\leq\dim G_{k}-2$ by 10.3.
By [42, Theorem 10.3] there is a closed reductive subgroup $H$ of $G$ defined over $\bar{k}$ and $g\in G(\kappa(\bar{\eta}))$ such that $H_{x}=g(H_{\bar{\eta}})g^{-1}$. Since $H$ is defined by finitely many polynomials with coefficients in $\bar{k}$, $H$ is defined over a finite extension of $k$. The point $x^{\prime}:=g(x)\in\overline{X}^{\mathrm{gen}}_{H}(\kappa(\bar{\eta}))$ maps to $\eta$ in $\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$, thus $\eta$ lies in $\overline{X}^{\mathrm{git}}_{HG,\overline{\rho}^{\mathrm{ss}}}$. ∎
@@ 13.3. Bounding the dimension of the space
In this subsection we carry out the induction to prove (DIM). It follows from [48, Theorem 9.3 (3)] that (DIM) holds if $H^{0}$ is a torus. Thus we may assume that $\dim H\geq 3$.
@@
Let $\overline{{V}}_{P,ij}$ be the preimage of $\overline{U}_{P,ij}:=\overline{U}_{LG,\overline{\rho}^{\mathrm{ss}}}\cap\overline{X}^{\mathrm{git}}_{H_{ij}L,\overline{\rho}^{\mathrm{ss}}}$ in $\overline{X}^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and let $\overline{Z}_{P,ij}$ be its closure. If part (a) holds then Corollary 13.6 gives us
@@
Lemma 9.11 implies that $\dim N_{k}\geq 1$, where $N$ is the unipotent radical of any R-parabolic with R-Levi $L$. It follows from Lemma 2.16, (73), (74) and (75) that $\dim\overline{Z}>\dim\overline{X}^{\mathrm{git}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ and thus $\dim\overline{X}^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}=\dim\overline{Z}$, which implies the last equality in (70). ∎
@@
If $\pi_{1}(G^{\prime})$ is étale or $x\in X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ then $(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})_{x}=(\operatorname{Lie}G^{\prime})_{x}$. It follows from Proposition 9.7 and Lemma 13.8 that $x\in V^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ if and only if $\rho_{x}$ is absolutely $G$-irreducible and $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))=0$. Since $x$ is closed in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$ its residue field is a local field by [6, Lemmas 3.17, 3.18]. Local Tate duality gives us an isomorphism $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))\cong H^{2}(\Gamma_{F},\operatorname{ad}^{0}\rho_{x})^{*}$, which implies the assertion. ∎
@@
where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map and $r=\dim G^{\prime}_{k}([F:\mathbb{Q}_{p}]+1)$. It is proved in [48, Corollary 9.5] that $X^{\mathrm{gen}}_{G/G^{\prime},\varphi\circ\overline{\rho}^{\mathrm{ss}}}$ is $\mathcal{O}$-flat of relative dimension $\dim(G/G^{\prime})_{k}([F:\mathbb{Q}_{p}]+1)$. It follows from 5.16 that $R^{\square}_{G/G^{\prime},\varphi\circ\rho_{x}}$ is flat over the coefficient ring $\Lambda$ for $\kappa(x)$ of relative dimension $\dim(G/G^{\prime})_{k}([F:\mathbb{Q}_{p}]+1)$. Thus $R^{\square}_{G,\rho_{x}}$ is flat over $\Lambda$ of relative dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$.
@@
For the purpose of establishing (DIM), after passing to a finite extension of $\mathcal{O}$ we may assume, that $G^{0}$ is split, see 5.9. Let $\mu:=Z(G^{0})\cap G^{\prime}=Z(G^{\prime})$, $G_{1}:=G/\mu$ and let $\varphi_{1}:G\to G_{1}$ be the projection map. Since $Z(G^{0})$ is a split torus, $\mu$ is a finite diagonalisable group scheme and we may decompose $\mu$ canonically into a $p$-part and a prime-to-$p$ part, which are both normal in $G$. Let $x$ be a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}}$. 3.12 and 3.16 imply that $\dim\overline{R}^{\square}_{G,\rho_{x}}=\dim\overline{R}^{\square}_{G_{1},\varphi_{1}\circ\rho_{x}}$.